OpenStudy (anonymous):

1. Graph each system of constraints. 3x+y≤7 x+2y≤9 x≥0,y≥0 2. Name all vertices. 3. Then find the values of x and y that maximize or minimize the objective function. 4. Find the maximum value.

OpenStudy (anonymous):

\[1.~x \ge 0,y \ge 0 ~\implies~the~graph~lies~\in~first~quadrant.\] 2. draw lines 3x+y=7,x+2y=9 3. find the points of intersection of these lines 4. find the points of intersection of these lines with x-axis by plugging y=0 and with y-axis by plugging x=0 5.put x=0,y=0 in \[3 x+y \le 7 ~and~x+2 y \le 9\] and then find the possible region. 6. find~the ~value~at~the~vertices.

OpenStudy (anonymous):

Thats right